OpenStudy (anonymous):

A winter jacket is marked down 75% from its original price. According to the new price tag, it now costs $52.50. What was the original price?

OpenStudy (campbell_st):

mark down of 75% means you pay 25% then 25% = $52.50 4 x 25% = 4 x $52.50 100% = 210

OpenStudy (anonymous):

\[\LARGE x(1-75\%)=52.50\] \[\LARGE x(1-\cfrac{75}{100 })=52.50\] \[\LARGE x-\cfrac{75x}{100 }=52.50\] \[\LARGE \frac{100x}{100}-\cfrac{75x}{100 }=52.50\] \[\LARGE \frac{25x}{100}=52.50\] \[\LARGE \frac{x}{4}=52.50\] can you do it now ? :)